OpenStudy (anonymous):

Choose the equation of a line that is perpendicular to the given line and that passes through the given point. 4x – 12y = 2; (10, –1) y = 3x + 29 y = –x + 29 y = –3x + 29 y = –x + 7 Choose the equation of a line that is perpendicular to the given line and that passes through the given point. 4x – 12y = 2; (10, –1) y = 3x + 29 y = –x + 29 y = –3x + 29 y = –x + 7 @Mathematics

OpenStudy (anonymous):

Find the slope y=mx+b y=1/3x - 1/6 so slope is 1/3 plug 1/3 for m and (10,1) for x, y into point-slope or slope-intercept

OpenStudy (anonymous):

(10,-1)

OpenStudy (anonymous):

use -3 for new slope

OpenStudy (anonymous):

you meen to use the neg recip so use -3 y=-3(x-10) -1

OpenStudy (anonymous):

y=-3x +29

OpenStudy (anonymous):

that is correct!